OpenStudy (aacollins223):

Object A attracts object B with a gravitational force of 10 newtons from a given distance. If the distance between the two objects is doubled, what is the changed force of attraction between them?

OpenStudy (irishboy123):

\(F = \dfrac{Gm_A~ m_B}{d^2} = 10\) \(F\color{red}{'} = \dfrac{Gm_A~ m_B}{(\color{red}{2d})^2} = \dfrac{1}{4} \dfrac{Gm_A~ m_B}{d^2}= \dfrac{1}{4} \cdot 10= 2.5\)